# Building reliable sim driving agents by scaling self-play

### v2

@@ 1 Introduction
Simulation agents are a core part of safely developing and testing systems that interact with humans, such as autonomous vehicles (AVs). In the context of self-driving, these agents, also referred to as road user behavior models, serve two primary purposes: establishing benchmarks for AV behavior (Engström et al., 2024), and representing other road users in simulators to enable statistical safety testing in both nominal and rare, long-tail scenarios (Corso et al., 2021; Montali et al., 2024). While each use case brings particular requirements, reliability is an important one that they share.
@@
How can we build sim agents that are close enough [^1] Here, close enough is emphasized because what constitutes an acceptable model of human behavior depends highly on the use case. to reality while maximizing designer specifications i.e. reliability? One approach relies on generative models, which have shown remarkable progress in producing diverse, human-like behaviors through imitation learning from demonstrations (Xu et al., 2023; Philion et al., 2024; Huang et al., 2024). However, whether they meet the reliability standards of a fully automated AV development pipeline is uncertain. This is highlighted by the top-performing models in the Waymo Open Sim Agent Challenge (Montali et al., 2024, WOSAC), a well-known benchmark for realistic nominal road user behavior. While state-of-the-art models in the 2024 challenge closely replicate logged human trajectories and achieve high scores on various distributional metrics, they still fall short in critical areas. Ground-truth human trajectories in the dataset rarely or never involve collisions or off-road movements, yet the top submissions (1st and 2nd place) frequently display such unintended behaviors. Specifically, simulated agents collide with others in 5–6% of scenarios and go off-road in 6–12% of cases (Zhou et al., 2024; Huang et al., 2024, BehaviorGPT, VBD).
@@
Figure 1: Overview of approach. Left: We define several criteria to guide the learning of simulation agents through rewards. The reward function is a weighted combination of these criteria: $r(o^{i}_{t})=\sum_{i}c_{i}\cdot\mathbb{I}[\text{criteria}_{i}]$. Here, we focus on achieving goal-directed nominal sim agent behavior—ensuring agents stay on the road and avoid collisions while navigating to a target position. Right: Over 24 hours on a single GPU, we iterate through 10,000 scenarios (green curve) from the Waymo Open Motion Dataset in GPUDrive (Kazemkhani et al., 2024), reaching near-perfect performance (blue curve, reliability) on the defined criteria after 2 billion agent steps by self-play PPO. The example scenarios illustrate agent behavior at different stages of training. Initially, agents display random behavior and frequently collide with each other and the road edges (marked in orange and red), but their behavior becomes streamlined over many iterations.
@@
This limits the scalability of AV evaluation and development, especially as generative models are increasingly used to create rare safety-critical scenarios underrepresented in real-world data (Mahjourian et al., 2024). When trajectories deviate unpredictably, researchers or engineers must find out: is the observed outcome a signal or an artifact of simulator noise? For instance, if 1 in 10 scenarios reflects unintended behavior, distinguishing meaningful failures from artifacts becomes a time-consuming task. As such, making sim agents more reliable seems a key pillar to further scale AV evaluation and development.
@@
The question becomes: how can we close this reliability gap in state-of-the-art sim agents? Assuming we can precisely define what the agent should adhere to (e.g. stay on the road), there is reason to believe that self-play reinforcement learning (RL) could be a piece of the puzzle. Evidence from a broad body of recent literature on games shows that self-play RL, combined with well-defined criteria (e.g. maximize score X) can produce agents capable of perfect, superhuman, gameplay in the large compute and data regime (Silver et al., 2018; Bakhtin et al., 2023; Berner et al., 2019).
@@
To investigate these questions, we train agents via self-play using a semi-realistic human perception framework in a data-driven simulator (Kazemkhani et al., 2024). We evaluate performance across thousands of scenarios from the Waymo Open Motion Dataset (Ettinger et al., 2021). Our key finding is that self-play PPO scales effectively with on-policy data and compute. After sufficient training, models generalize well to 10,000 unseen test traffic scenarios, virtually closing the train-test gap.
@@ 2.1 Dataset and simulator
We conduct our experiments in GPUDrive, a data-driven, multi-agent, GPU-accelerated simulator (Kazemkhani et al., 2024). GPUDrive contains $K=$ 160,147 real-world traffic scenarios from the Waymo Open Motion Dataset (Ettinger et al., 2021, WOMD). Each scenario $k\in K$ comprises a static road graph, $R_{k}$, and a time series of joint logged human trajectories:
@@ 2.2.2 Metrics
The Collided and Off-road metrics align with the Waymo Open Sim Agent Challenge and Waymax (Montali et al., 2024; Gulino et al., 2024). Specifically, Collided is part of the “object interaction metrics” category and the off-road events are part of the “map-based metrics” category. Under the assumption that human road users have near zero collision and off-road events, we can meaningfully compare our scores to the top submissions (Huang et al., 2024; Zhou et al., 2024)  [^2] Technically, WOSAC frames this as a distribution-matching problem: metrics are first computed as event counts, which are then compared to the distribution of log replay trajectories across several rollouts..
The Goal achieved metric is not directly reported in WOSAC, making it less comparable. The most similar metric is the Route Progress Ratio used in Waymax (Gulino et al., 2024), which measures how far an agent travels along the logged trajectory. However, since our focus is not on mimicking logged trajectories but on precisely reaching a particular goal, a binary metric is, in our case, a more meaningful indicator of performance. However, reaching the goal roughly corresponds to a Route Progress Ratio of $100\%$.
@@ 2.3 State and observation space
This section outlines the design choices and parameterization of the observation $\mathbf{o}_{t}^{i}$ for agent $i$ at time $t$. We make these choices to reflect semi-realistic limits on human perception. The observation encodes the agent’s partial view of the scenario state $s_{t}$, capturing the information necessary for decision-making. In this work, we model the RL problem as a Partially Observed Stochastic Game (Hansen et al., 2004, POSG), where agents make simultaneous decisions under partial observability. We further make the following design choices for our agents:
@@
We reduce the full road graph, which consists of up to 10,000 sparsely distributed road points, in dimension for computational efficiency. To reduce the number of points corresponding to straight lines, we run the polyline reduction threshold of the polyline decimation algorithm (Visvalingam & Whyatt, 2017) in GPUDrive to 0.1 which roughly cuts the number of points by a factor of 10. We also cap the maximum visible road points at 200, selecting 200 points from those in the view radius in a random order if there are more than $200$ points, creating a sparse view of the local road graph. Empirical results show this is sufficient for agents to navigate toward goals without going off the road or causing collisions.
@@ 2.4 Action space and dynamics model
the conditional pmf $\pi_{\theta}$ describes the behavior under the assumption that $\mathbf{o}_{t}^{i}$ takes a fixed set of values. The total joint action space contains $7\times 13=91$ actions. With these actions, agents are stepped in the simulator using an Ackermann bicycle model (Rajamani, 2011).
@@ 2.5 Reward function
Here, $\mathbb{I}[.]$ is an indicator function that equals 1 if the condition is true and 0 otherwise. We assign weights $w_{\text{Goal Achieved}}=1.0$, $w_{\text{Collided}}=0.75$, and $w_{\text{Offroad}}=0.75$. An agent achieves the goal position when it reaches within a 2-meter radius of the target $(x,y)$. Once an agent reaches its goal, we remove it from the scene. This latter choice is made as it is ill-defined what an agent should do after it reaches its goal.
@@ 2.7 Models
We use a neural network with an encoder and a shared embedding, as illustrated in Figure 3. The flat observation vector is first decomposed into three modalities: the dense ego state, the sparse road graph, and the sparse partner observations. Each modality is processed independently. Inspired by the late fusion approach in Wayformer (Nayakanti et al., 2023), we then concatenate the outputs, apply max pooling, and pass the result through a shared embedding. This hidden embedding is fed into separate actor and critic heads, each implemented as a single feedforward layer. The model only has $\approx 50,000$ trainable parameters.
@@ 2.8 Training
In each scenario, we control up to $N=64$ agents using a shared, decentralized policy $\pi_{\theta}$. Actions are independently sampled from the policy based on the ego views of each agent $i$ during every step in the rollout: $\mathbf{a}_{t}^{i}\sim\pi_{\theta}(\cdot\mid\mathbf{o}_{t}^{i})$. We train agents using Proximal Policy Optimization (Schulman et al., 2017, PPO) using batches of $S=800$ distinct scenarios, with the set of training scenarios uniformly resampled every 2 million steps. Initially, agents exhibit random behavior and crash frequently. Over time, the agents’ behavior becomes more streamlined, creating smooth trajectories with high rates of reaching the goals.
@@ 3 Related work
Self-play RL (Samuel, 1959; Tesauro et al., 1995) has been a core ingredient in creating effective agents across a wide range of complex games. Notable examples include superhuman gameplay in two-player zero-sum games like Chess and Go (Silver et al., 2018), expert human-level play in Stratego (Perolat et al., 2022) and Starcraft (Vinyals et al., 2019), as well many-player games that require some level of cooperation like Diplomacy (Bakhtin et al., 2023) and Gran Turismo (Wurman et al., 2022). These successes have demonstrated the effectiveness of self-play, particularly in the large-data, large-compute regime. However, the majority of its successes are in variants of zero-sum games whereas driving tasks are likely general-sum and feature many-agent interaction.
@@
Reinforcement learning has been explored for the design of autonomous driving agents, though state-of-the-art agents are currently far below the human rate of between $800000$ km per police-reported traffic crash in the United States (Stewart et al., 2023) or as much as $1$ crash per $24800$ km in more challenging domains such as San Francisco (Flannagan et al., 2023). These agents are frequently trained in simulators built atop large open-source driving datasets (Gulino et al., 2024; Vinitsky et al., 2022; Kazemkhani et al., 2024) such as Waymo Open Motion (Ettinger et al., 2021, WOMD), (Caesar et al., 2020, NuScenes), (Zheng et al., 2024, ONE-Drive) though there are also procedurally generated (Li et al., 2022) and non-data-driven simulators (Dosovitskiy et al., 2017). These datasets collectively add up to tens of thousands of hours of available data and are often used to train RL agents in log-replay mode, a setting in which only one agent is learning and the remainder are either replaying human trajectories or executed hand-coded policies. The complexity of scaling RL in these settings has led to the creation of batched simulators (Kazemkhani et al., 2024, GPUDrive), (Gulino et al., 2024, Waymax) whose high throughput helps ameliorate issues of sample complexity. Many works have explored ways to use these simulators to learn high-quality reinforcement learning agents through RL including uses of self-play (Peng et al., 2021; Vinitsky et al., 2022; Zhang et al., 2022, 2023, 2024). Our work is mostly distinct from these by the scale of training and a significantly lower crash and off-road rate than has previously been observed.
@@ 4.1 Scaling with data
We investigate whether agents with a partial view of the environment can solve all scenarios in the Waymo Open Motion Dataset. Our results show that nearly all scenarios can be solved successfully. After $1$ billion training steps, agents achieve a goal-reaching rate of $99.84$%, a collision rate of $0.38$%, and an off-road rate of $0.26$% on the training dataset. Furthermore, as depicted in Figure 5, zooming in on the final four hours of training suggests that metrics exhibit a continued, albeit gradual, improvement, indicating that performance can be further increased with additional training. This training run took $24$ hours on a single NVIDIA A100 GPU. [^3] These metrics are computed in alignment with the way they are defined in WOSAC, but it should be noted that this is an over-optimistic metric as it includes many agents that simply need to remain in place as they are initialized right next to their goals. This initialization mode can be reproduced in the simulator by setting init_mode = all_objects. Excluding such agents, the performance metrics are: $99.40$% goal-reaching rate, $0.5$% collision rate, and $0.6$% offroad rate. The latter initialization mode, referred to as init_mode = all_non_trivial, only controls agents that must drive more than 2 meters before reaching their goal and is used during training.
@@ 5 Discussion
Our findings suggest that self-play RL scales effectively with available data (Section 4.1), achieving state-of-the-art performance on the Waymo Open Motion Dataset (WOMD) with no generalization gap. To the best of our knowledge, this is the first demonstration of this level of performance on WOMD. Compared to state-of-the-art supervised models, such as VBD (Huang et al., 2024) and BehaviorGPT (Zhou et al., 2024), our approach reduces collision and off-road rates by at least 15 $\times$.
@@ 5.2 Concluding thoughts
We demonstrated the potential of scaling self-play to develop agents that can be precisely controlled to meet specific criteria in autonomous driving. While not explored in this paper, we anticipate that our findings extend to other domains such as neuroscience, where agent-based modeling is gaining momentum (Aldarondo et al., 2024; Johnson-Yu et al., ; Castro et al., 2025). In neuroscience, researchers are increasingly using physics-based simulators to create digital twins of animals, enabling cost-effective and controlled experimentation. For these agents to be useful models of animal behavior, reliability and robustness appear essential. A rodent foraging model, for example, should not exhibit free movement. We hope our work contributes to the improvement of agent-based modeling, helping to enhance controllability and robustness across different domains.

### v3

@@ 1 Introduction
Simulation agents are a core part of safely developing and testing systems that interact with humans, such as autonomous vehicles (AVs). In the context of self-driving, these agents, also referred to as road user behavior models, serve two primary purposes: establishing benchmarks for AV behavior (engstrom2024modeling), and representing other road users in simulators to enable statistical safety testing in both nominal and rare, long-tail scenarios (corso2021survey; montali2024waymo). While each use case brings particular requirements, reliability is an important one that they share.
@@
How can we build sim agents that are close enough [^1] Here, close enough is emphasized because what constitutes an acceptable model of human behavior depends highly on the use case. to reality while maximizing designer specifications i.e. reliability? One approach relies on generative models, which have shown remarkable progress in producing diverse, human-like behaviors through imitation learning from demonstrations (xu2023bits; DBLP:conf/iclr/PhilionPF24; huang2024versatile). However, whether they meet the reliability standards of a fully automated AV development pipeline is uncertain. This is highlighted by the top-performing models in the Waymo Open Sim Agent Challenge (montali2024waymo, WOSAC), a well-known benchmark for realistic nominal road user behavior. While state-of-the-art models in the 2024 challenge closely replicate logged human trajectories and achieve high scores on various distributional metrics, they still fall short in critical areas. Ground-truth human trajectories in the dataset rarely or never involve collisions or off-road movements, yet the top submissions (1st and 2nd place) frequently display such unintended behaviors. Specifically, simulated agents collide with others in 5–6% of scenarios and go off-road in 6–12% of cases (zhou2024behaviorgpt; huang2024versatile, BehaviorGPT, VBD).
@@
Figure 1: Overview of approach. Left: We define several criteria to guide the learning of simulation agents through rewards. The reward function is a weighted combination of these criteria: $r(o^{i}_{t})=\sum_{i}c_{i}\cdot\mathbb{I}[\text{criteria}_{i}]$. Here, we focus on achieving goal-directed nominal sim agent behavior—ensuring agents stay on the road and avoid collisions while navigating to a target position. Right: Over 24 hours on a single GPU, we iterate through 10,000 scenarios (green curve) from the Waymo Open Motion Dataset in GPUDrive (kazemkhani2024gpudrive), reaching near-perfect performance (blue curve, reliability) on the defined criteria after 2 billion agent steps by self-play PPO. The example scenarios illustrate agent behavior at different stages of training. Initially, agents display random behavior and frequently collide with each other and the road edges (marked in orange and red), but their behavior becomes streamlined over many iterations.
@@
This limits the scalability of AV evaluation and development, especially as generative models are increasingly used to create rare safety-critical scenarios underrepresented in real-world data (mahjourian2024unigen). When trajectories deviate unpredictably, researchers or engineers must find out: is the observed outcome a signal or an artifact of simulator noise? For instance, if 1 in 10 scenarios reflects unintended behavior, distinguishing meaningful failures from artifacts becomes a time-consuming task. As such, making sim agents more reliable seems a key pillar to further scale AV evaluation and development.
@@
The question becomes: how can we close this reliability gap in state-of-the-art sim agents? Assuming we can precisely define what the agent should adhere to (e.g. stay on the road), there is reason to believe that self-play reinforcement learning (RL) could be a piece of the puzzle. Evidence from a broad body of recent literature on games shows that self-play RL, combined with well-defined criteria (e.g. maximize score X) can produce agents capable of perfect, superhuman, gameplay in the large compute and data regime (silver2018general; DBLP:conf/iclr/Bakhtin0LGJFMB23; openai5).
@@
To investigate these questions, we train agents via self-play using a semi-realistic human perception framework in a data-driven simulator (kazemkhani2024gpudrive). We evaluate performance across thousands of scenarios from the Waymo Open Motion Dataset (ettinger2021large). Our key finding is that self-play PPO scales effectively with on-policy data and compute. After sufficient training, models generalize well to 10,000 unseen test traffic scenarios, virtually closing the train-test gap.
@@ 2.1 Dataset and simulator
We conduct our experiments in GPUDrive, a data-driven, multi-agent, GPU-accelerated simulator (kazemkhani2024gpudrive). GPUDrive contains $K=$ 160,147 real-world traffic scenarios from the Waymo Open Motion Dataset (ettinger2021large, WOMD). Each scenario $k\in K$ comprises a static road graph, $R_{k}$, and a time series of joint logged human trajectories:
@@ 2.2.2 Metrics
The Collided and Off-road metrics align with the Waymo Open Sim Agent Challenge and Waymax (montali2024waymo; gulino2024waymax). Specifically, Collided is part of the “object interaction metrics” category and the off-road events are part of the “map-based metrics” category. Under the assumption that human road users have near zero collision and off-road events, we can meaningfully compare our scores to the top submissions (huang2024versatile; zhou2024behaviorgpt)  [^2] Technically, WOSAC frames this as a distribution-matching problem: metrics are first computed as event counts, which are then compared to the distribution of log replay trajectories across several rollouts..
The Goal achieved metric is not directly reported in WOSAC, making it less comparable. The most similar metric is the Route Progress Ratio used in Waymax (gulino2024waymax), which measures how far an agent travels along the logged trajectory. However, since our focus is not on mimicking logged trajectories but on precisely reaching a particular goal, a binary metric is, in our case, a more meaningful indicator of performance. However, reaching the goal roughly corresponds to a Route Progress Ratio of $100\%$.
@@ 2.3 State and observation space
This section outlines the design choices and parameterization of the observation $\mathbf{o}_{t}^{i}$ for agent $i$ at time $t$. We make these choices to reflect semi-realistic limits on human perception. The observation encodes the agent’s partial view of the scenario state $s_{t}$, capturing the information necessary for decision-making. In this work, we model the RL problem as a Partially Observed Stochastic Game (posg04, POSG), where agents make simultaneous decisions under partial observability. We further make the following design choices for our agents:
@@
We reduce the full road graph, which consists of up to 10,000 sparsely distributed road points, in dimension for computational efficiency. To reduce the number of points corresponding to straight lines, we run the polyline reduction threshold of the polyline decimation algorithm (visvalingam2017line) in GPUDrive to 0.1 which roughly cuts the number of points by a factor of 10. We also cap the maximum visible road points at 200, selecting 200 points from those in the view radius in a random order if there are more than $200$ points, creating a sparse view of the local road graph. Empirical results show this is sufficient for agents to navigate toward goals without going off the road or causing collisions.
@@ 2.4 Action space and dynamics model
the conditional pmf $\pi_{\theta}$ describes the behavior under the assumption that $\mathbf{o}_{t}^{i}$ takes a fixed set of values. The total joint action space contains $7\times 13=91$ actions. With these actions, agents are stepped in the simulator using an Ackermann bicycle model (rajamani2011vehicle).
@@ 2.7 Models
We use a neural network with an encoder and a shared embedding, as illustrated in Figure 3. The flat observation vector is first decomposed into three modalities: the dense ego state, the sparse road graph, and the sparse partner observations. Each modality is processed independently. Inspired by the late fusion approach in Wayformer (DBLP:conf/icra/NayakantiAZGRS23), we then concatenate the outputs, apply max pooling, and pass the result through a shared embedding. This hidden embedding is fed into separate actor and critic heads, each implemented as a single feedforward layer. The model only has $\approx 50,000$ trainable parameters.
@@ 2.8 Training
In each scenario, we control up to $N=64$ agents using a shared, decentralized policy $\pi_{\theta}$. Actions are independently sampled from the policy based on the ego views of each agent $i$ during every step in the rollout: $\mathbf{a}_{t}^{i}\sim\pi_{\theta}(\cdot\mid\mathbf{o}_{t}^{i})$. We train agents using Proximal Policy Optimization (ppoSchulman17, PPO) using batches of $S=800$ distinct scenarios, with the set of training scenarios uniformly resampled every 2 million steps. Initially, agents exhibit random behavior and crash frequently. Over time, the agents’ behavior becomes more streamlined, creating smooth trajectories with high rates of reaching the goals.
@@ 3 Related work
Self-play RL (5392560; tesauro1995temporal) has been a core ingredient in creating effective agents across a wide range of complex games. Notable examples include superhuman gameplay in two-player zero-sum games like Chess and Go (silver2018general), expert human-level play in Stratego (perolat2022mastering) and Starcraft (starcraft), as well many-player games that require some level of cooperation like Diplomacy (DBLP:conf/iclr/Bakhtin0LGJFMB23) and Gran Turismo (gtsophy). These successes have demonstrated the effectiveness of self-play, particularly in the large-data, large-compute regime. However, the majority of its successes are in variants of zero-sum games whereas driving tasks are likely general-sum and feature many-agent interaction.
@@
Reinforcement learning has been explored for the design of autonomous driving agents, though state-of-the-art agents are currently far below the human rate of between $800000$ km per police-reported traffic crash in the United States (stewart2023overview) or as much as $1$ crash per $24800$ km in more challenging domains such as San Francisco (flannagan2023establishing). These agents are frequently trained in simulators built atop large open-source driving datasets (gulino2024waymax; nocturne; kazemkhani2024gpudrive) such as Waymo Open Motion (ettinger2021large, WOMD), (caesar2020nuscenes, NuScenes), (onedrive, ONE-Drive) though there are also procedurally generated (li2022metadrive) and non-data-driven simulators (carla17). These datasets collectively add up to tens of thousands of hours of available data and are often used to train RL agents in log-replay mode, a setting in which only one agent is learning and the remainder are either replaying human trajectories or executed hand-coded policies. The complexity of scaling RL in these settings has led to the creation of batched simulators (kazemkhani2024gpudrive, GPUDrive), (gulino2024waymax, Waymax) whose high throughput helps ameliorate issues of sample complexity. Many works have explored ways to use these simulators to learn high-quality reinforcement learning agents through RL including uses of self-play (copo; nocturne; closed_loop_driving; closed_loop_v2; aspDrive). Our work is mostly distinct from these by the scale of training and a significantly lower crash and off-road rate than has previously been observed.
@@ 4.1 Scaling with data
We investigate whether agents with a partial view of the environment can solve all scenarios in the Waymo Open Motion Dataset. Our results show that nearly all scenarios can be solved successfully. After $2$ billion training steps, agents achieve a goal-reaching rate of $99.84$%, a collision rate of $0.38$%, and an off-road rate of $0.26$% on the training dataset. Furthermore, as depicted in Figure 5, zooming in on the final four hours of training suggests that metrics exhibit a continued, albeit gradual, improvement, indicating that performance can be further increased with additional training. This training run took $24$ hours on a single NVIDIA A100 GPU. [^3] These metrics are computed in alignment with the way they are defined in WOSAC, but it should be noted that this is an over-optimistic metric as it includes many agents that simply need to remain in place as they are initialized right next to their goals. This initialization mode can be reproduced in the simulator by setting init_mode = all_objects. Excluding such agents, the performance metrics are: $99.40$% goal-reaching rate, $0.5$% collision rate, and $0.6$% offroad rate. The latter initialization mode, referred to as init_mode = all_non_trivial, only controls agents that must drive more than 2 meters before reaching their goal and is used during training.
@@ 5 Discussion
Our findings suggest that self-play RL scales effectively with available data (Section 4.1), achieving state-of-the-art performance on the Waymo Open Motion Dataset (WOMD) with no generalization gap. To the best of our knowledge, this is the first demonstration of this level of performance on WOMD. Compared to state-of-the-art supervised models, such as VBD (huang2024versatile) and BehaviorGPT (zhou2024behaviorgpt), our approach reduces collision and off-road rates by at least 15 $\times$.
@@ 5.2 Concluding thoughts
We demonstrated the potential of scaling self-play to develop agents that can be precisely controlled to meet specific criteria in autonomous driving. While not explored in this paper, we anticipate that our findings extend to other domains such as neuroscience, where agent-based modeling is gaining momentum (aldarondo2024virtual; johnson2024understanding; castro2025discovering). In neuroscience, researchers are increasingly using physics-based simulators to create digital twins of animals, enabling cost-effective and controlled experimentation. For these agents to be useful models of animal behavior, reliability and robustness appear essential. A rodent foraging model, for example, should not exhibit free movement. We hope our work contributes to the improvement of agent-based modeling, helping to enhance controllability and robustness across different domains.